# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ IV Convergence Analysis
$\leq$ $\sum_{i\in\mathcal{N}}\Bigg[-\frac{1}{2}\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{U}_{i}}^{2}-\frac{\kappa_{1}-1}{2}\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{H}_{i}^{T}\mathbf{H}_{i}}^{2}$
- $\frac{c(\kappa_{1}-1)}{2}\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}-\frac{\rho}{4}\|\Delta\mathbf{w}_{i}^{t+1}\|^{2}$
- $\frac{c(\kappa_{1}-6(2N_{\max}+1))}{2}\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}$
- $\frac{c\kappa_{1}-6(N_{\max}+2)}{2c}\|\mathbf{H}_{i}\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{D}_{i}\Delta\mathbf{w}_{i}^{t}\|^{2}$
- $\frac{c\kappa_{1}-6(1+2c)(N_{\max}+2)}{2}\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2}$
- $\left(\frac{c\kappa_{2}\tilde{\tau}_{\min\quad}}{6N_{\text{sum}}n(c+1)^{2}}-\frac{(2N_{\max}+1)c\kappa_{1}}{2}\right)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}$
@@
where $N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}$ represents the total number of neighboring nodes in the sensor network, $d_{\max}:=\max\left(\{d_{i,j}\mid i\in\mathcal{N},j\in\mathcal{N}_{i}\}\cup\{r_{i}\mid i\in\mathcal{N}\}\right)$ denotes the maximum distance measurement, and $\tilde{\tau}_{\min}:=\min\left\{5(c+1)^{2}N_{i}^{2}+(3c^{2}+4c+3)N_{i},\;i\in\mathcal{N}\right\}.$
@@
$\kappa_{1}\geq 6(2N_{\max}+2)\left(1+\frac{1}{c}\right).$ (40)
@@
$\kappa_{2}\geq\frac{3N_{\text{sum}}n(c+1)^{2}(2N_{\max}+1)\kappa_{1}}{\tilde{\tau}_{\min\quad}}.$ (41)
@@
For each node $i\in\mathcal{N}$, suppose the sequence $\{(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t},\bm{\lambda}_{i}^{t})\}_{t\geq 1}$ is generated by Algorithm 1. Then, the following conclusions hold:
@@
$\mathbf{0}$ $\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\partial\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i}^{t}),\quad\forall i\in\mathcal{N}$ (43b)
@@
Moreover, $(\mathbf{x}^{t},\mathbf{y}_{1}^{t},\mathbf{w}^{t})$ is a critical point of problem (10), and $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ is a critical point of the original problem (2).
@@
[…]
@@
A tuple $(\mathbf{z},\mathbf{w},\bm{\lambda})$ is called an $\epsilon$-solution of problem (24) if it satisfies $\|\mathbf{A}_{i}\mathbf{z}_{i}\|<\epsilon$ and there exist vectors $\bm{\eta}_{1}\in\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\mathbf{A}_{i}\bm{\lambda}_{i}+\partial\delta_{\mathcal{X}\times\mathcal{Y}}(\mathbf{z}_{i})$, $\bm{\eta}_{2}\in\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\partial\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i})$ such that $\|\bm{\eta}_{1}\|\leq\epsilon$ and $\|\bm{\eta}_{2}\|\leq\epsilon$.
@@
See Appendix G. ∎
V Numerical Experiments
We conduct experiments on a two-dimensional localization problem to validate the proposed SP-ADMM-JCNL algorithm. To examine the algorithm’s performance and applicability under different settings, the experimental evaluation is organized into two parts: (i) performance assessment on a synthetic sensor network with controllable topology and noise levels, which enables flexible investigation of the algorithm’s behavior under different parameter settings and network configurations; and (ii) comparison with a two-stage separate localization strategy (described in Algorithm 2) on a real-world benchmark dataset, where sensor and target positions are estimated in sequential steps.
To quantitatively assess the results, we separately report the root mean squared error (RMSE) for sensor self-localization and target localization. This distinction allows for a clearer understanding of the respective performance in both cooperative and non-cooperative components of the joint localization task.
@@ V Numerical Experiments
    $\text{RMSE}_{\text{sensor}}(t)=\sqrt{\frac{\sum\limits_{i\in(\mathcal{N}/\mathcal{A})}\|\mathbf{x}_{i}^{t}-\mathbf{x}_{i}\|^{2}}{N-m}},$ (44)
@@
  • The target localization RMSE at iteration $t$ is given by:
    $\text{RMSE}_{\text{target}}(t)=\sqrt{\frac{\sum\limits_{i\in\mathcal{N}}\|\mathbf{y}_{i}^{t}-\mathbf{y}\|^{2}}{N}},$ (45)
    where $\mathbf{y}_{i}^{t}$ represents the estimated target position by sensor $i$ and $\mathbf{y}$ is the true target location.
@@
V-A Synthetic Network Evaluation
We begin by constructing a synthetic sensor network to evaluate the convergence behavior and localization accuracy of the proposed SP-ADMM-JCNL algorithm in a controlled environment. Specifically, a total of 100 agent nodes, 8 anchors, and one target are randomly deployed within a square region $0,1]\times[0,1]$. The communication range is set to $C_{\text{range}}=0.3$, which results in a moderately connected network topology with sparse local neighborhoods. This level of connectivity reflects realistic conditions often encountered in low-power wireless sensor deployments. The corresponding network topology used in our experiments is shown in Fig. 2.
@@ V-A Synthetic Network Evaluation
Following [34], we consider two types of measurement noise to simulate different sensing conditions. The first type is additive white Gaussian noise (AWGN), where the measurement noise is modeled as a zero-mean Gaussian variable with fixed variance, i.e., $\sigma_{i,j}^{2}=\sigma_{\text{add}}^{2}$. This noise model captures the baseline level of measurement uncertainty commonly assumed in localization literature. The second type is range-dependent Gaussian noise, where the variance increases proportionally with the square of the Euclidean distance between node pairs. Specifically, the noise is modeled as zero-mean Gaussian with $\sigma_{i,j}^{2}=\sigma_{\text{add}}^{2}\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$, where $\|\mathbf{p}_{i}-\mathbf{p}_{j}\|$ denotes the true distance between nodes $i$ and $j$. This setting reflects more realistic scenarios where measurement errors grow with distance. In this part, we set $\sigma_{\text{add}}=0.02$. These two noise types are tested separately in order to evaluate the robustness and adaptability of the proposed algorithm under varying noise characteristics.
For the synthetic network experiments, the penalty parameters in the augmented Lagrangian are set to $\rho=c=0.11$ and the initial dual variable is set as $\mathbf{u}^{0}=\mathbf{0}$. Initial estimates are drawn from a uniform distribution: $z_{0}\sim\text{Unif}(-1,1)^{(8\mathcal{E}+2N)}$, where $\mathcal{E}$ denotes the set of communication links and $N$ is the number of agents.
@@
We first examine the convergence behavior of the proposed algorithm. Figures 3(a) and 3(c) respectively show the RMSE of the sensors and the target as a function of the number of iterations when additive white Gaussian noise (AWGN) and distance-dependent Gaussian noise are used. To further examine the convergence of the SP-ADMM algorithm, we plot the optimality gap of problem (24) versus the number of iterations in Figures 3(b) and 3(d) . It can be clearly observed that the performance gap of the proposed SP-ADMM algorithm decreases as the number of iterations increases. This result validates Theorem 1, which states that the iterative sequence $\{(\mathbf{z}^{t},\mathbf{u}^{t},\bm{\lambda}^{t})\}$ generated by Algorithm 1 converges to a KKT critical point of problem (24), with a convergence rate of $\mathcal{O}(1/T)$. Figure 4 shows the joint localization results of sensors and the target obtained by the SP-ADMM-JCNL algorithm in the scenario depicted in Figure 3. The first row presents results under additive white Gaussian noise (AWGN), while the second row corresponds to range-dependent Gaussian noise model. In both settings, the proposed method is able to accurately recover the sensor and target positions. The estimated locations closely match the ground truth, demonstrating the robustness of the algorithm across different noise types. These visual results align well with the quantitative RMSE evaluations discussed earlier.
@@
Figure 4: Localization results with synthetic network. Measurement noise: AWGN (first row) and range-dependent Gaussian noise(second row). Anchors:  ; true agent positions: +; true target positions: +; estimated agent positions: $\circ$; estimated target positions: $\circ$.
V-B Benchmark Network
In this section, we evaluate the performance of the proposed SP-ADMM-JCNL algorithm against a baseline that separately addresses cooperative and non-cooperative localization tasks. For clarity, we refer to this baseline as SP-ADMM-SCNL. The corresponding two-stage procedure is summarized in Algorithm 2, where sensor self-localization is performed first using anchor-sensor range measurements, followed by target localization based on the estimated sensor positions.
We consider a two-dimensional localization problem and utilize the publicly available benchmark dataset test10-500, which can be accessed at http://www.stanford.edu/yyye/. In this dataset, sensor nodes are randomly deployed within the unit square $0.5,0.5]\times[-0.5,0.5]$, and the ground-truth positions of both sensors and anchors are provided. This setup allows for consistent and reproducible evaluation across different methods. Since the dataset does not include any target node information, we manually assign the true position of a passive target at the origin, $(0,0)$, and compute range measurements from each sensor to the target accordingly. For both JCNL and SCNL, the same noisy measurements are used as input to ensure fair comparison. All experiments are repeated over multiple Monte Carlo trials, and the average RMSE for sensor and target localization is reported to assess accuracy and robustness.
@@ V-B Benchmark Network
Input: Anchor positions $\{\mathbf{a}_{k}\}_{k=1}^{m}$, node-to-node distances, parameters $c$, $\rho$, initial values $\mathbf{z}_{i}^{0}$, $\mathbf{w}_{i}^{0}$, $\bm{\lambda}_{i}^{0}=\mathbf{0}$.
@@
Stage 2: Target Localization
@@
  Treat the estimated $\{\hat{\mathbf{x}}_{i}\}$ as virtual anchors (i.e., set $m=N$) then apply SP-ADMM-JCNL to estimate the target position $\hat{\mathbf{y}}$.
@@
The measurement noise follows a range-dependent Gaussian model and we set $\sigma_{\text{add}}=0.02$. Initial estimates are drawn from a uniform distribution: $z_{0}\sim\text{Unif}(-1,1)^{(8\mathcal{E}+2N)}$, where $\mathcal{E}$ denotes the set of communication links and $N$ is the number of agents. The penalty parameters in the augmented Lagrangian are set to $c=\rho=0.18$.For the proposed joint localization method (SP-ADMM-JCNL), we set the iteration cap to $T=4000$. For the two-stage separate localization strategy (SP-ADMM-SCNL), we allocate $T_{1}=2000$ iterations for the sensor self-localization stage and $T_{2}=2000$ iterations for the subsequent target localization stage, keeping the total iteration budget consistent. To ensure statistical reliability and account for randomness in initialization and noise, we conduct 100 independent Monte Carlo trials. The final results are presented in terms of average RMSE values over these runs.
@@
Figure 5 illustrates the localization results for both sensor self-localization and target localization under the two different processing strategies: the separate approach (SP-ADMM-SCNL) and the proposed joint approach (SP-ADMM-JCNL). For the separate method, the average computation time was approximately 0.0626 seconds for the first stage (cooperative localization of sensors) and 0.0863 seconds for the second stage (non-cooperative localization of the target), resulting in a total runtime of around 0.1489 seconds. In contrast, the joint method required an average computation time of about 0.0955 seconds to complete both tasks simultaneously, which is slightly lower than the cumulative cost of the two-stage approach.
The efficiency gain arises from the ability of the joint method to perform simultaneous updates of sensor and target positions, which eliminates the sequential dependency inherent in the separate strategy. This structural advantage leads to faster convergence, especially in the early iterations, as can be seen in the RMSE curves for target localization. This property is especially advantageous in real-time or dynamic environments, where delays in target estimation could affect subsequent system responses. Moreover, the joint formulation ensures coherent updates between sensor and target updates throughout the optimization process, reducing the risk of error accumulation that may arise from treating the two stages separately. These results support the practical value of the proposed JCNL model in scenarios where real-time inference and coherence across multiple localization tasks are essential.
VI Conclusion

### v2

@@ 4 Convergence Analysis
$\leq$ $\sum_{i\in\mathcal{N}}\Bigg[-\frac{(\kappa_{1}-1)}{2}(c\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}+\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{H}_{i}^{T}\mathbf{H}_{i}}^{2})$
- $\frac{1}{2}\|\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{U}_{i}}^{2}-\frac{c(\kappa_{1}-6L_{2})}{2}\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t+1}\|_{\mathbf{A}_{i}^{T}\mathbf{A}_{i}}^{2}$
- $\frac{\rho}{4}\|\Delta\mathbf{w}_{i}^{t+1}\|^{2}-\frac{c\kappa_{1}-6L_{1}}{2c}\|\mathbf{H}_{i}\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{D}_{i}\Delta\mathbf{w}_{i}^{t}\|^{2}$
- $\left(\frac{c\kappa_{2}\tilde{\tau}_{\min\quad}}{2(2M+1)N_{\text{sum}}n(c+1)^{2}}-\frac{L_{2}c\kappa_{1}}{2}\right)\|\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|^{2}$
- $\frac{c\kappa_{1}-6(1+c)L_{1}}{2}\|\Delta\tilde{\mathbf{z}}_{i}^{t+1}-\Delta\mathbf{z}_{i}^{t}\|_{\mathbf{B}_{i}^{T}\mathbf{B}_{i}}^{2}$
@@
where $L_{1}=N_{\max}+M_{\max}+1,L_{2}=2N_{\max}+1$, with $N_{\max}:=\max\{N_{i}:i\in\mathcal{N}\}$, $M_{\max}:=\max\{M_{i}:i\in\mathcal{M}\}$. While $N_{\text{sum}}:=\sum_{i\in\mathcal{N}}N_{i}$ represents the total number of neighboring nodes in the sensor network, $d_{\max}:=\max\left(\{d_{i,j}\mid i\in\mathcal{N},j\in\mathcal{N}_{i}\}\cup\{r_{i,{\color[rgb]{0,0,0}k}}\mid i\in\mathcal{N},{\color[rgb]{0,0,0}k\in\mathcal{M}_{i}\}}\right)$ denotes the maximum distance measurement, and $\tilde{\tau}_{\min}:=\min\left\{{\color[rgb]{0,0,0}(c+1)^{2}[(6M+1)N_{i}^{2}+N_{i}]},\;i\in\mathcal{N}\right\}.$
@@
$\kappa_{1}\geq 6(2N_{\max}+{\color[rgb]{0,0,0}M_{\max}}+1)\left(1+\frac{1}{c}\right).$ (39)
@@
$\kappa_{2}\geq\frac{{\color[rgb]{0,0,0}(2M+1)}N_{\text{sum}}n(c+1)^{2}(2N_{\max}+1)\kappa_{1}}{\tilde{\tau}_{\min\quad}}.$ (40)
@@
Consider the sequence $\{(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t},\bm{\lambda}_{i}^{t})\}_{t\geq 1},\forall i\in\mathcal{N}$ produced by Algorithm 1. Then:
@@
$\mathbf{0}$ $\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{t},\mathbf{w}_{i}^{t})+\partial\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i}^{t}),\quad\forall i\in\mathcal{N}$ (42b)
@@
Moreover, $(\mathbf{x}^{t},\mathbf{y}_{1}^{t},\mathbf{w}^{t})$ is a critical point of problem (9), and $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ is a critical point of the original problem (2), where $\mathbf{y}_{1}^{t}=\text{vec}(\mathbf{y}_{1,k},k\in\mathcal{M})$.
@@
Since $\mathbf{z}\in\mathcal{Y}$, if $(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ satisfies condition (42c), then it follows that the constraint $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}=\mathbf{y}_{j,k},\forall(i,j)\in\mathcal{E},{\color[rgb]{0,0,0}k\in\mathcal{M}}$ holds. Given that the sensor network is connected, this implies the consistency of the entire sensor network’s estimation of the target position, i.e., for any $i\in\mathcal{N},k\in\mathcal{M}$, we have $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}=\mathbf{y}_{1,k}^{t}$. Here, $\mathbf{y}_{1}^{t}$ can be replaced with any $\mathbf{y}_{i}^{t},i\in\mathcal{N}$.
@@
A tuple $(\mathbf{z},\mathbf{w},\bm{\lambda})$ is called an $\epsilon$-solution of problem (23) if it satisfies $\|\mathbf{A}_{i}\mathbf{z}_{i}\|<\epsilon$ and there exist vectors $\bm{\eta}_{1}\in\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\mathbf{A}_{i}\bm{\lambda}_{i}+\partial\delta_{\mathcal{X}\times\mathcal{Y}}(\mathbf{z}_{i})$, $\bm{\eta}_{2}\in\nabla_{\mathbf{z}_{i}}G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\partial\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i})$ such that $\|\bm{\eta}_{1}\|\leq\epsilon$ and $\|\bm{\eta}_{2}\|\leq\epsilon$.
@@
See Appendix A.7. ∎
5 Numerical Experiments
We conduct experiments on a two-dimensional localization problem to validate the proposed SP-ADMM-JCNL algorithm. To examine the algorithm’s performance and applicability under different settings, the experimental evaluation is organized into two parts: (i) performance assessment on a synthetic sensor network with controllable topology and noise levels, which enables flexible investigation of the algorithm’s behavior under different parameter settings and network configurations; and comparison with a two-stage separate localization strategy (described in Algorithm 2), where sensor and target positions are estimated in sequential steps. (ii) analysis of the impact of different parameter configurations on the performance and convergence of JCNL.
To quantitatively assess the results, we separately report the root mean squared error (RMSE) for sensor self-localization and targets localization. This distinction allows for a clearer understanding of the respective performance in both cooperative and non-cooperative components of the joint localization task.
@@ 5 Numerical Experiments
    $\text{RMSE}_{\text{sensor}}(t)=\sqrt{\frac{\sum\limits_{i\in(\mathcal{N}/\mathcal{A})}\|\mathbf{x}_{i}^{t}-\mathbf{x}_{i}\|^{2}}{N-{\color[rgb]{0,0,0}A}}},$ (43)
@@
  • The targets localization RMSE at iteration $t$ is given by:
    $\text{RMSE}_{\text{{\color[rgb]{0,0,0}target}}}(t)=\sqrt{\frac{\sum\limits_{{\color[rgb]{0,0,0}k\in\mathcal{M}}}\sum\limits_{i\in\mathcal{N}}\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{y}_{k}\|^{2}}{MN}},$ (44)
    where $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}$ represents the estimated target position by sensor $i$ and $\mathbf{y}$ is the true target location.
@@
5.1 Synthetic Network Evaluation
We begin by constructing a synthetic sensor network to evaluate the convergence behavior and localization accuracy of the proposed SP-ADMM-JCNL algorithm in a controlled environment. Specifically, a total of 50 agents, 8 anchors, and four targets are randomly deployed within a square region $0,1]\times[0,1]$. The communication range is set to $C_{\text{range}}=0.3$, which results in a moderately connected network topology with sparse local neighborhoods. This level of connectivity reflects realistic conditions often encountered in low-power wireless sensor deployments. The corresponding network topology used in our experiments is shown in Fig. 2.
@@ 5.1 Synthetic Network Evaluation
Following [31], we consider two types of measurement noise to simulate different sensing conditions. The first type is additive white Gaussian noise (AWGN), where the measurement noise is modeled as a zero-mean Gaussian variable with fixed variance, i.e., $\sigma_{i,j}^{2}=\sigma_{\text{add}}^{2}$. This noise model captures the baseline level of measurement uncertainty commonly assumed in localization literature. The second type is range-dependent Gaussian noise, where the variance increases proportionally with the square of the Euclidean distance between sensor pairs. Specifically, the noise is modeled as zero-mean Gaussian with $\sigma_{i,j}^{2}=\sigma_{\text{add}}^{2}\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$, where $\|\mathbf{p}_{i}-\mathbf{p}_{j}\|$ denotes the true distance between sensors $i$ and $j$. This setting reflects more realistic scenarios where measurement errors grow with distance. In this part, we set $\sigma_{\text{add}}^{2}=0.02$. These two noise types are tested separately in order to evaluate the robustness and adaptability of the proposed algorithm under varying noise characteristics.
For the synthetic network experiments, the penalty parameters in the augmented Lagrangian are set to $\rho=c=0.26$ and the initial dual variable is set as $\mathbf{u}^{0}=\mathbf{0}$. Initial estimates are drawn from a uniform distribution: $\mathbf{z}_{0}\sim\text{Unif}(-1,1)^{(M+1)(4|\mathcal{E}|+2N)}$, where $\mathcal{E}$ denotes the set of communication links and $N$ is the number of sensors.
@@
We first examine the convergence behavior of the proposed algorithm. Figures 3(a) and 3(c) respectively show the RMSE of the sensors and the targets as a function of the number of iterations when additive white Gaussian noise (AWGN) and distance-dependent Gaussian noise are used. To further examine the convergence of the SP-ADMM algorithm, we plot the optimality gap of problem (23) versus the number of iterations in Figures 3(b) and 3(d) . It can be clearly observed that the performance gap of the proposed SP-ADMM algorithm decreases as the number of iterations increases. This result validates Theorem 1, which states that the iterative sequence $\{(\mathbf{z}^{t},\mathbf{u}^{t},\bm{\lambda}^{t})\}$ generated by Algorithm 1 converges to a KKT critical point of problem (23), with a convergence rate of $\mathcal{O}(1/T)$. Figure 4 shows the joint localization results of sensors and targets obtained by the SP-ADMM-JCNL algorithm in the scenario depicted in Figure 2. The first row presents results under additive white Gaussian noise (AWGN), while the second row corresponds to range-dependent Gaussian noise model. In both settings, the proposed method is able to accurately recover the sensor and target positions. The estimated locations closely match the ground truth, demonstrating the robustness of the algorithm across different noise types. These visual results align well with the quantitative RMSE evaluations discussed earlier.
@@
Figure 4: Localization results with synthetic network. Measurement noise: AWGN and range-dependent Gaussian noise. Anchors:  ; true agent positions: +; true target positions: +; estimated agent positions: $\circ$; estimated target positions: $\circ$.
Moreover, we evaluate the performance of the proposed SP-ADMM-JCNL algorithm against a baseline that separately addresses cooperative and non-cooperative localization tasks. For clarity, we refer to this baseline as SP-ADMM-SCNL. The corresponding two-stage procedure is summarized in Algorithm 2, where sensor self-localization is performed first using anchor-sensor range measurements, followed by targets localization based on the estimated sensor positions.
We consider a two-dimensional localization problem with a randomly generated network comprising $100$ sensors, $12$ anchors, and $10$ targets uniformly distributed in $0,1]\times[0,1]$. Range measurements are corrupted by additive white Gaussian noise with variance $\sigma_{\text{add}}^{2}=0.01$, and the communication range is set to $C_{range}=0.25$. For both JCNL and SCNL, identical noisy measurements are used as input. All experiments are repeated over $100$ Monte Carlo trials, and the average RMSE for sensor and target localization is reported. For both JCNL and SCNL, the same noisy measurements are used as input to ensure fair comparison. All experiments are repeated over multiple Monte Carlo trials, and the average RMSE for sensors and targets localization is reported to assess accuracy and robustness.
@@
Input: Anchor positions $\{\mathbf{a}_{l}\}_{l=1}^{{\color[rgb]{0,0,0}A}}$, sensor -to-sensor distances, parameters $c$, $\rho$, initial values $\mathbf{z}_{i}^{0}$, $\mathbf{w}_{i}^{0}$, $\bm{\lambda}_{i}^{0}=\mathbf{0}$.
@@
Stage 2: targets localization
@@
  Treat the estimated $\{\hat{\mathbf{x}}_{i}\}$ as virtual anchors (i.e., set ${\color[rgb]{0,0,0}A}=N$) then apply SP-ADMM-JCNL to estimate the targets position $\{\hat{\mathbf{y}}_{k}\}$.
@@
Initial estimates are drawn from a uniform distribution: $\mathbf{z}_{0}\sim\text{Unif}(-1,1)^{(M+1)(4|\mathcal{E}|+2N)}$, where $\mathcal{E}$ denotes the set of communication links and $N$ is the number of sensors. The penalty parameters in the augmented Lagrangian are set to $c=\rho=0.18$. For the proposed joint localization method (SP-ADMM-JCNL), we set the iteration cap to $T=5000$. For the two-stage separate localization strategy (SP-ADMM-SCNL), we allocate $T_{1}=2000$ iterations for the sensor self-localization stage and $T_{2}=3000$ iterations for the subsequent targets localization stage, keeping the total iteration budget consistent. To ensure statistical reliability and account for randomness in initialization and noise, we conduct 100 independent Monte Carlo trials. The final results are presented in terms of average RMSE values over these runs.
@@
Figure 5 illustrates the localization results for both sensor self-localization and targets localization under the two different processing strategies: the separate approach (SP-ADMM-SCNL) and the proposed joint approach (SP-ADMM-JCNL). For the separate method, the average computation time was approximately 0.0258 seconds for the first stage (cooperative localization of sensors) and 0.0863 seconds for the second stage (non-cooperative localization of the targets), resulting in a total runtime of around 0.1643 seconds. In contrast, the joint method required an average computation time of about 0.1768 seconds to complete both tasks simultaneously, which is slightly lower than the cumulative cost of the two-stage approach. The joint method exhibits faster convergence in the initial iterations, particularly in estimating the target positions. This improvement is crucial in real-time applications. Moreover, unlike the separate approach, the joint strategy avoids target estimation delays by updating sensor and target locations simultaneously.
5.2 Parameter Selection and Sensitivity Analysis
This section evaluates the localization performance of the proposed SP-ADMM-JCNL algorithm under different choices of penalty parameters $c$, $\rho$ and initial values $\mathbf{u}^{0}$. The goal is to investigate the sensitivity of JCNL to these algorithmic parameters and to provide empirical guidance for selecting suitable values in practice. These results can help users reduce the effort required for parameter tuning when applying JCNL to new localization tasks.
5.2.1 Parameter Selection for $\rho$ and $c$
According to Theorem 2, if the parameter $\rho$ is selected to satisfy conditions (39)–(41) for a given $c$, the SP-ADMM algorithm is guaranteed to converge. We test the algorithm on the network configuration depicted in Fig. 2, which consists of $N=50$ sensors, $m=8$ anchors, and $M=4$ targets where nodes are uniformly distributed over a $0,1]\times[0,1]$ square region.
Figure 6: RMSE curves achieved by our proposed methods with different penalty parameters $\rho$ under the benchmark network (N = 50, m = 8, M=4) with Range-dependent noise. $RMSE_{sensor}$(left column); $RMSE_{target}$(right column).
Figure 6 respectively shows the RMSE performance of sensor self-localization and target localization achieved by JCNL under different values of $\rho$ with a fixed $c=0.18$. The results indicate that the choice of $\rho$ and $c$ affects both the convergence speed and final localization accuracy of the proposed algorithm. Specifically, an excessively large $\rho$ leads to slower convergence and reduced localization accuracy, as evidenced by the red and light blue curves in Figure 6. It can be observed that when $\rho$ lies within the range $[0.09c,2.0c]$, the performance of JCNL remains relatively stable. And figure 7 respectively presents the RMSE performance for sensor self-localization and target localization obtained under different values of $c$ with a fixed $\rho=0.18$. It is observed that both excessively large and excessively small values of $c$ result in slower convergence and degraded localization accuracy, as shown by the red, light blue, dark blue and orange curves in Figure 7. In particular, when $c$ is selected within the range $0.5\rho,2.0\rho]$, the algorithm demonstrates stable and favorable performance.
Figure 7: RMSE curves achieved by our proposed methods with different penalty parameters $c$ under the benchmark network (N = 50, m = 8, M=4) with Range-dependent noise. $RMSE_{sensor}$(left column); $RMSE_{target}$(right column).
5.2.2 Localization Performance With $u_{0}$ Selection:
Figure 8 illustrates the influence of the initialization parameter $u_{0}$ on sensor self-localization and target localization, respectively. It is observed that the performance exhibits only minor fluctuations with respect to the choice of $u_{0}$. These results suggest that the proposed SP-ADMM-JCNL algorithm is relatively insensitive to the initialization of the parameter $u_{0}$.
Figure 8: RMSE curves achieved by our proposed methods with different penalty parameters $\rho$ under the benchmark network (N = 50, m = 8, M=4) with Range-dependent noise. $RMSE_{sensor}$(left column); $RMSE_{target}$(right column).
6 Conclusion